Project Cash Flows Your company, Dawgs “R” Us, is evaluating a new project involving the purchase of a new oven to bake your hotdog buns. If purchased, the new oven will replace your existing oven, which was purchased seven years ago for a total installed price of $1 million. You have been depreciating the old oven on a straight-line basis over its expected life of 15 years to an ending book value of $250,000, even though you expect it to be worthless at the end of that 15-year period. The new oven will cost $2 million and will fall into the MACRS five-year depreciation class life. If you purchase the new oven, you expect it to last for eight years. At the end of those eight years, you expect to be able to sell it for $100,000. (Note that both of the ovens, old and new, therefore have an effective remaining life of eight years at the time of your analysis.) If you do purchase the new oven, you estimate that you can sell the old one for its current book value at the same time. The advantages of the new oven are twofold: Not only do you expect it to reduce the before-tax costs on your current baking operations by $75,000 per year, but you will also be able to produce new types of buns. The sales of the new buns are expected to bring your company $200,000 per year throughout the eight-year life of the new oven, while associated costs of the new buns are only expected to be $80,000 per year. Since the new oven will allow you to sell these new products, you anticipate that NWC will have to increase immediately by $20,000 upon purchase of the new oven. It will then remain at that increased level throughout the life of the new oven to sustain the new, higher level of operations. Your company uses a required rate of return of 12 percent for such projects, and your incremental tax rate is 34 percent. What will be the total cash flows for this project? Note: MACRS 5 yr is as follows; 20%, 32%, 19.20%, 11.52%, 11.52%, 5.76%
Answer #1

Answer:

Existing Oven:

Installed 7 years ago at cost = $1,000,000

Annual depreciation = (1000000 - 250000) / 15 = $50,000

Book value now = 1000000 - 7 * 50000 = $650,000

Hence sale value now = $650,000

As required cash flows from year 0 to Year 8 are calculated and given below:

1 2 3 4 5 Year New Oven cost Sale proceeds of old oven Increase in NWC ($2,000,000) $650,000 ($20,000) Saving in before tax c